# QKD security proofs for decoy-state BB84: protocol variations, proof techniques, gaps and limitations

### v1

@@ 6.4 Entropy accumulation based proofs
There are a variety of accumulation theorems, such as the entropy accumulation theorem (EAT) [97, 102], generalized entropy accumulation theorem (GEAT) [46, 103] and generalized Rényi entropy accumulation theorem (GREAT) [103], which we discuss in greater detail below. Currently, most of these results are focused on fixed-length protocols. Any of these theorems can be utilized for QKD security proofs, as long as the QKD protocol can be abstractly represented as a sequential process satisfying the conditions of the relevant theorem. An example of such a sequential process can be seen in Fig. 3. However, in general one needs to be careful to verify that a sequential process indeed correctly represents a QKD protocol and satisfies the assumptions of the desired accumulation theorem.
@@
Abstractly, entropy accumulation theorems obtain a lower bound on the $n$-round entropy (smooth-min and/or Rényi) of a sequential process in terms of a single-round quantity of the form [^52] The correction term here depends on the probability of the event $\Omega_{\mathrm{acc}}$.
@@
where $\mathbf{S}$ are the secret registers containing the pre-amplification string and $\mathbf{E}$ is all of Eve’s side information except the communication for error correction, error verification and the choice of the hash function for privacy amplification. Here $n$ is the total number of rounds in the QKD protocol, and $h$ is a single-round quantity that typically depends on a construction called the min-tradeoff function, which we will discuss in more detail below. In particular, $h$ is found by minimizing the min-tradeoff function over the acceptance set.
Furthermore, as described in Section 4.5, having a valid lower bound on the entropy of the state conditioned on the protocol accepting is the essential ingredient for a security proof. The EAT, GEAT and GREAT provide powerful tools to achieve such lower bounds. Additionally, they do not require a security proof against IID attacks; they enable one to directly bound the entropy, comparable to the entropic uncertainty relation (Section 6.1).
6.4.1 Dupuis et al., 2020 [97]
This work proves the first version of the EAT for a sequential process (generating the pre-amplification string) that satisfies a technical requirement called the Markov condition. The Markov condition restricts the theorem to protocols that make only “weakly dependent announcements” [104] and limits it to being applied mainly to entanglement-based protocols. A weakly dependent announcement requires, for example, that a detection event depends only on the total photon number, but not the state. All of the results based on the original EAT perform the reduction to a single-round quantity by calculating the so-called min-tradeoff function, which is an affine lower bound on the key rate in terms of announcements in the test rounds, e.g. the error rate in the $X$-basis. In Ref. [104], a numerical framework for constructing min-tradeoff functions of entanglement-based protocols in the EAT-framework was presented, and then used to compute key rates for entanglement-based protocols.
@@
Ref. [46] extended the above result to a generalized entropy accumulation theorem (GEAT), which considers sequential processes satisfying a different technical requirement called the non-signalling condition (instead of the Markov condition). This allowed it to be applied to prepare-and-measure protocols; however, only under a specific restriction. Namely, it requires that one must assume that while Eve can jointly attack all signals, she only interacts with one signal at a time, i.e. applies the attack in sequence. We shall refer to this assumption as the “sequential assumption”. Satisfying this assumption in practical implementations imposes strong limits on the repetition rate at which the protocol can be run. Ref. [106] uses the GEAT to compute key rates for the B92 protocol, and describes the application of the GEAT to decoy-state calculations, without performing explicit calculations.
@@
Note that in general, the task of numerically finding the min-tradeoff function with the best finite-size performance can be very difficult. Thus, most works so far have focused on applying EAT- or GEAT-based security proofs to simple device-independent QKD protocols or qubit protocols, with Ref. [31] described below being the only work using it to explicitly compute key rates for decoy-state protocols. Ref. [103] improves the GEAT by obtaining tighter finite-size key rates, and avoiding explicit min-tradeoff function calculations. This leads to a significantly simpler framework, but it has not been applied to decoy-state protocols yet.
6.4.3 Kamin, Arqand et al. 2024 [31]
@@
This work applies the GEAT [46, 106] to the qubit BB84 and decoy-state BB84 protocols. While the usage of the GEAT requires the sequential assumption discussed above, this requirement can be removed with the results in Refs. [107, 108] (or Ref. [109], currently in preparation). Furthermore, its security proof only applies to fixed-length protocols.
@@
One technical contribution of this work was that it developed several techniques to optimize the finite-size performance of the GEAT, for instance by optimizing the choice of the min-tradeoff function using techniques somewhat similar to Ref. [104]. Moreover, a tighter formulation for decoy-state protocols is derived, which even in the asymptotic regime should result in higher key rates than the approaches in Ref. [110]. Finally, in Ref. [31, App. B] it was shown that one can incorporate the flag-state squasher with the GEAT analysis, which is especially relevant for optical detection setups.
@@ Acknowledgments
The writing of this document greatly benefited from interactions with the wider community, which includes numerous insights, discussions, and feedback, over many years. However, any errors or omissions are entirely our own responsibility. We would like to thank Takaya Matsuura for teaching the details of the phase error correction proof approach. We would like to thank Masato Koashi for further explaining the applications of the phase error correction proof approach, and the source-replacement scheme as applied to this proof technique. We thank Fred Fung for explaining the universal squasher to us. We thank Toyohiro Tsurumaru, Victor Zapatero, Margarida Pereira and Guillermo Currás-Lorenzo for helpful discussions about the EUR and phase error correction techniques, and the similarities between them. We also thank Ian George, Jie Lin and Peter Brown for helpful discussions on the EAT and its improved versions.

### v2

@@ 6.4 Entropy accumulation based proofs
There are a variety of accumulation theorems, such as the entropy accumulation theorem (EAT) [31, 30], generalized entropy accumulation theorem (GEAT) [67, 68] generalized Rényi entropy accumulation theorem [3], and marginal-constrained entropy accumulation theorems (MEAT) [46, 32, 4], which we discuss in greater detail below. Currently, most of these results are focused on fixed-length protocols. Any of these theorems can be utilized for QKD security proofs, as long as the QKD protocol can be abstractly represented as a sequential process satisfying the conditions of the relevant theorem. An example of such a sequential process can be seen in Fig. 3. However, in general one needs to be careful to verify that a sequential process indeed correctly represents a QKD protocol and satisfies the assumptions of the desired accumulation theorem.
@@
Abstractly, entropy accumulation theorems obtain a lower bound on the $n$-round entropy (smooth-min and/or Rényi) of a sequential process in terms of a single-round quantity of the form [^57] The correction term here depends on the probability of the event $\Omega_{\mathrm{acc}}$. Technically, the sequential process outputs a string different from $\mathbf{S}$, which goes through a discarding procedure, before resulting in the pre-amplification string $\mathbf{S}$. However, this discrepancy can be resolved as discussed earlier in Section 5.6.2.
@@
where $\mathbf{S}$ are the secret registers containing the pre-amplification string and $\mathbf{E}$ is all of Eve’s side information except the communication for error correction, error verification and the choice of the hash function for privacy amplification. Here $n$ is the total number of rounds in the QKD protocol, and $h$ is a value that can be computed in terms of a minimization problem only involving single rounds of the protocol.
6.4.1 Ref. [31]
This work proves the first version of the EAT for a sequential process (generating the pre-amplification string) that satisfies a technical requirement called the Markov condition. The Markov condition restricts the theorem to protocols that make only “weakly dependent announcements” [37] and limits it to being applied mainly to entanglement-based protocols. A weakly dependent announcement requires, for example, that a detection event depends only on the total photon number, but not the state. All of the results based on the original EAT perform the reduction to the single-round quantity $h$ by calculating the so-called min-tradeoff function, which is an affine lower bound on the key rate in terms of announcements in the test rounds, e.g. the error rate in the $X$-basis. In Ref. [37], a numerical framework for constructing min-tradeoff functions of entanglement-based protocols in the EAT-framework was presented, and then used to compute key rates for entanglement-based protocols.
@@
Ref. [67] extended the above result to a generalized entropy accumulation theorem (GEAT), which considers sequential processes satisfying a different technical requirement called the non-signalling condition (instead of the Markov condition). This allowed it to be applied to prepare-and-measure protocols; however, only under a specific restriction. Namely, it requires that one must assume that while Eve can jointly attack all signals, she only interacts with one signal at a time, i.e. applies the attack in sequence. We shall refer to this assumption as the “sequential assumption”. Satisfying this assumption in practical implementations imposes strong limits on the repetition rate at which the protocol can be run [90, Section 5.1]. Ref. [68] uses the GEAT to compute key rates for the B92 protocol, and describes the application of the GEAT to decoy-state calculations, without performing explicit calculations.
@@
Note that in general, the task of numerically finding the min-tradeoff function (required for GEAT calculations) with the best finite-size performance can be very difficult. Thus, most works so far have focused on applying EAT- or GEAT-based security proofs to simple device-independent QKD protocols or qubit protocols, with Ref. [49] described below being the only work using it to explicitly compute key rates for decoy-state protocols. Ref. [3] improves the GEAT by obtaining tighter finite-size key rates, and avoiding explicit min-tradeoff function calculations, which leads to a significantly simpler framework.
6.4.3 Ref. [49]
@@
This work applies the GEAT [67, 68] to the qubit BB84 and decoy-state BB84 protocols. Technically, the usage of the GEAT in this work requires the sequential assumption discussed above, though it was noted in that work that this requirement can be removed by instead using the “marginal-constrained” versions [4, Corollary 4.1] we describe next (Section 6.4.4). Furthermore, its security proof only applies to fixed-length protocols.
@@
One technical contribution of this work was that it developed several techniques to optimize the finite-size performance of the GEAT, for instance by optimizing the choice of the min-tradeoff function using techniques somewhat similar to Ref. [37]. Moreover, a tighter formulation for decoy-state protocols is derived, which even in the asymptotic regime should result in higher key rates than the approaches in Ref. [51]. Finally, in Ref. [49, App. B] it was shown that one can incorporate the flag-state squasher with the GEAT analysis, which is especially relevant for optical detection setups.
6.4.4 Refs. [32, 4, 50]
Refs. [32, 4] derived [^58] The results in those two works are generalizations of a result in [46] that has thus far only been presented via conference talks, but we note that they both provide standalone proofs that do not depend on that result. a new marginal-constrained entropy accumulation theorem (MEAT) that is particularly well-suited for the analysis of prepare-and-measure protocols. This result enables analysis of both fixed-length and variable-length protocols, while also offering a simpler calculation compared to the GEAT in [67]. Building on this, Ref. [50] develops a numerical framework for computing key rates for generic decoy-state protocols using the MEAT technique, and applies it to compute key rates for the decoy-state BB84 and the decoy-state 4-6 protocol. It also includes the analysis of certain device imperfections. We note that while the MEAT version in [4] can in principle be used for handling on-the-fly announcements, Ref. [50] does not consider this scenario.
7 Summary of properties of different proof techniques
As seen in the previous sections, security proofs for QKD protocols are complex, and rely on a wide variety of mathematical tools and arguments. Given that there exist a variety of proof approaches, in this section we will briefly assess their respective strengths and weaknesses. In doing so, we lay the foundation to make an informed choice of proof techniques for the decoy-state BB84 protocol. We stress that this assessment of the various proof techniques reflects the authors’ perspective. A choice of the proof technique inevitably involves trade-offs that different individuals may judge differently. Moreover, the characteristics of proof techniques continue to evolve as this remains an active area of research.
In comparing the proof techniques, we consider the following aspects:
  1. 1.
    Performance: The ability of the proof technique to yield tight key rates in the finite-size regime.
  2. 2.
    Robust against device imperfections: The ability of the proof technique to be modified to address imperfections in the physical hardware, including both IID and non-IID deviations, in contrast to the technique requiring idealized assumptions.
  3. 3.
    Scrutiny and Peer Review: The extent to which the proof approach has been extensively analyzed, scrutinized, and undergone peer review, and whether it is widely accepted in the community.
  4. 4.
    Modularity: The extent to which proof techniques can be adapted to different protocol variations without requiring a complete overhaul of the proof.
  5. 5.
    Accessibility: The extent to which they are accessible to experts well-versed in quantum information theory
  6. 6.
    Protocol Choices: The range of protocol variations that can be accommodated within the proof technique, such as active vs passive detection setups, advantage distillation, noisy preprocessing, on-the-fly announcements, etc.
  7. 7.
    Computability of Key Rates: The ease with which the technique can compute the key rates. In particular, the extent to which the proof technique requires difficult numerical optimizations, in contrast to using analytical formulae.
Table 3: Subjective comparison of proof techniques
Criterion Entropic uncertainty relation Phase error correction Postselection technique Entropy accumulation theorem
Performance Good key rates. Good key rates. Pessimistic key rates. Good key rates.
Robust against device imperfections Robust against (independent) device imperfections. Robust against (independent) device imperfections. Robust against (IID) device imperfections. Robust against (independent) device imperfections.
Scrutiny and Peer Review High degree of scrutiny. Oldest “modern” proof technique. High degree of scrutiny. High degree of scrutiny. Very recent work. Not much scrutiny .
Modularity The different extensions (e.g. for robustness against device imperfections, ability to make on-the-fly announcements, etc.) are not modular. Thus, it is not obvious that the extensions can be used simultaneously in the same proof. The different extensions (e.g. for robustness against device imperfections, ability to make on-the-fly announcements, etc.) are not modular. Thus, it is not obvious that the extensions can be used simultaneously in the same proof. Fairly modular and easy to apply. Fairly modular and easy to apply.
Accessibility Many proofs are accessible, and written in terms of well-defined mathematical statements. Not very accessible, many works are not written in terms of well-defined mathematical statements. Accessible, written in terms of well-defined mathematical statements. Accessible, written in terms of well-defined mathematical statements, although requires a good degree of familiarity with relevant technical machinery.
Protocol Choices Does not apply to passive protocols, can only handle certain on-the-fly announcements in idealized situations, requires certain classical postprocessing. It is unclear how noisy preprocessing and advantage distillation can be incorporated. Does not apply to passive protocols, can only handle certain on-the-fly announcements in idealized situations, requires certain classical postprocessing. Can only handle specific error-correction protocols. Requires a random permutation on the classical data. Cannot handle on-the-fly announcements. Cannot handle advantage distillation. The theoretical framework itself can handle on-the-fly announcements, but more work is required when reducing the key rate computation to a single-round convex optimization problem.
Computability of Key Rates Analytical key rates. Easy to compute. Analytical key rates. Easy to compute. Numerical key rates. Not too hard to compute. Numerical key rates. Harder to compute than the postselection technique.
7.1 Entropic uncertainty relations approach
  1. 1.
    Performance: This technique provides good key rates.
  2. 2.
    Robust against device imperfections: There exist some works that address hardware imperfections using this technique (however, we note that in some cases this requires a quite significant change to the overall structure of the proof). Additionally, due to the close similarity between this technique and phase error correction based methods, this approach can also leverage the numerous results that establish suitable phase error rate bounds in the presence of imperfections [25, 125, 42, 99, 73, 23, 96, 97, 123, 79, 113, 77]. Thus, we consider this proof technique to be robust against a wide range of device imperfections. However, we note that care must be taken when incorporating results from the phase error correction based works into the EUR framework [^59] Informally, this requires ensuring that the phase error rate bounds derived in the phase error correction framework are consistent with those needed in the EUR-based analysis.. Furthermore, not all existing works on device imperfections are inherently compatible with one another, thus care must be taken in combining these as well.
  3. 3.
    Scrutiny and Peer Review: The proofs are well-established, having undergone extensive peer review over many years [105, 106].
  4. 4.
    Modularity: Many proofs using this technique are somewhat modular, with well-stated theorems and lemmas, making them relatively structured and adaptable. However, this modularity is not necessarily retained when considering works that address imperfections and deviations from the ideal BB84 protocol, or protocol variations such as on-the-fly announcements.
  5. 5.
    Accessibility: We believe this proof technique to be accessible. Many works using this proof technique have clear theorem statements that can be understood by a reader well versed in quantum information [105].
  6. 6.
    Protocol choices:
    • Active vs passive protocols: Suitable bounds on the phase error rate have only been obtained for protocols with active detection setups. Thus, this technique (along with the phase error correction technique) can currently only be applied to active protocols. Some passive protocols (e.g. polarisation-encoded BB84 with perfectly balanced beam splitter) can be shown to be equivalent to active protocols, and thus this technique is applicable to them.
    • On-the-fly announcements: This technique can be applied to protocols that make certain kinds of announcements during the protocol [101]. However, currently no known method exists to combine this work with hardware imperfections.
    • Classical data processing: This technique must map double-click events to single-click events (in order to not have basis efficiency mismatch), which may prevent us from obtaining optimal key rates. It remains unclear how this proof technique can accommodate noisy preprocessing [^60] One might attempt to incorporate noisy preprocessing by just changing Alice’s POVM in Eq. 22, but this does not improve the $c_{q}$ parameter, leading to worse bounds on the final key rates after accounting for the increased error-correction cost. Thus, this straightforward attempt is not the correct way to obtain good key rate bounds with noisy preprocessing in the EUR approach. or advantage distillation.
  7. 7.
    Computability of Key Rates: This proof technique gives analytical formulae for key rates, which makes their computation straightforward.
7.2 Phase error correction approach
  1. 1.
    Performance: This technique provides good key rates.
  2. 2.
    Robust against device imperfections: The core technique here too involves idealized hardware assumptions, which have been relaxed with subsequent work. Incorporating device imperfections into this proof technique is a subject of active research [25, 125, 42, 99, 23, 96, 97, 123, 79, 113, 77], and there has been a large amount of progress on this front in recent years. Thus, we view this proof technique as robust to device imperfections. However, as with the EUR technique, not all works on imperfections are inherently compatible with each other, or with diverse protocol choices (see also Item 4).
  3. 3.
    Scrutiny and Peer Review: The proofs [56, 55, 57] are well-established, having undergone extensive peer review over many years.
  4. 4.
    Modularity: This proof technique is not stated in a modular fashion in the literature. Every modification seems to require a new proof, and it is not obvious if the different proof modifications fit with each other. For example, the current work on device imperfections is inherently incompatible with the work done to incorporate on-the-fly announcements into the proof [^61] Since both the EUR and phase error correction methods rely on many of the same works for phase error estimation, this observation is also applicable to the EUR proof technique.. Thus, care must be taken to ensure that security is proved for all protocol modifications simultaneously.
  5. 5.
    Accessibility: As a result, this proof technique is not very accessible. Additionally, papers often do not have clear theorem statements, further reducing the accessibility of this proof technique.
  6. 6.
    Protocol choices:
    • Active vs passive protocols: Similar to the EUR method, this technique cannot be applied to generic passive protocols due to the lack of a suitable proven bound on the phase error rate. Some passive protocols (e.g. polarisation-encoded BB84 with perfectly balanced beam splitter) can be shown to be equivalent to active protocols, and thus this technique is applicable to them.
    • On-the-fly announcements: Similar to the EUR, this technique can be applied to protocols that make certain kinds of announcements during the protocol [101] (see also Item 4).
    • Classical data processing: This technique must map double-click events to single-click events (in order to not have basis efficiency mismatch), which may prevent us from obtaining optimal key rates.
    • Furthermore, the basic proof technique is fairly restrictive in the kind of error correction that can be performed in the protocol (see Section 6.2 for a more detailed discussion). It is only very recently that noisy preprocessing and more complicated error correction protocols [28] have been incorporated into this proof technique, and this has not received as much scrutiny as the rest of the proof technique [^62] Additionally, this proof technique requires that the hash family used is linear and dual universal2, whereas the other three proof techniques use universal2 hashing. However, Toeplitz matrices and random binary matrices are both dual universal2 and universal2..
  7. 7.
    Computability of Key Rates: This proof technique gives analytical formulae for key rates, which makes their computation straightforward.
7.3 Postselection technique
  1. 1.
    Performance: This technique provides pessimistic key rates.
  2. 2.
    Robustness Against Device Imperfections: Since this proof technique reduces the security analysis to the IID setting, it is naturally compatible with a wide range of IID hardware imperfections [126, 120, 74, 76]. However, its ability to handle non-IID imperfections remains unclear. This is because ensuring permutation invariance of the protocol—a key requirement of this technique—is challenging, even when applying a random permutation to the classical data. As a result, extending this approach to non-IID imperfections is not straightforward.
  3. 3.
    Scrutiny and Peer Review: The proofs are well-established, having undergone extensive peer review over many years [22]. Critical gaps and extensions to practical scenarios were addressed in recent works [75].
  4. 4.
    Modularity: This proof technique is generally stated in well-defined and self-contained mathematical statements. More importantly, the theorem statements are fairly general and allow for generic protocols. Thus, it is modular — one needs to separately apply this technique for the lift to coherent attacks, and have protocol-specific proof steps for obtaining the IID security proof.
  5. 5.
    Accessibility: We believe this proof technique to be accessible to a reader well versed in quantum information, since it is stated in well-defined and self-contained mathematical statements.
  6. 6.
    Protocol choices:
    • Random permutation: This proof technique requires a random permutation to be applied to the classical data after the measurements (and IID behaviour of the hardware).
    • On-the-fly announcements: The technique fundamentally starts with a source-replaced state defined over the global protocol execution, and thus it is unclear how it can be modified to incorporate on-the-fly announcements.
    • Other protocol choices such as active vs passive basis choice, advantage distillation or noisy preprocessing are easily handled within this method.
  7. 7.
    Computability of Key Rates: This proof technique has moderate numerical requirements, since informally, one only has to perform key rate computations in the IID case (with suitably modified parameters) [38, 62, 120, 51].
7.4 Entropy accumulation approach
There are multiple variations of entropy accumulation, most prominently the original EAT [31], the subsequent GEAT [67], and very recently some “marginal-constrained” versions (which we shall broadly include under the term MEAT, for brevity) [46, 32, 4]. Out of these variations, the MEAT appears most suitable for analyzing PM protocols, since the EAT faces difficulties incorporating Alice’s marginal state constraints in the source-replacement technique, while applying the GEAT to a PM protocol currently requires it to satisfy a restrictive form of sequential structure that limits its repetition rate. Hence we shall focus on the MEAT for the points discussed below.
  1. 1.
    Performance: This technique provides good key rates.
  2. 2.
    Robustness against Device Imperfections: The reduction to single rounds in the MEAT only relies on fairly mild core assumptions (mostly regarding memory in the devices), and the analysis of single rounds would mainly rely on finding suitable source maps and squashing maps in the presence of device imperfections, which have been studied in various forms in the IID case. Hence we believe this approach should be robust to many types of imperfections; however, there are currently few works [50] that explicitly perform such an analysis.
  3. 3.
    Scrutiny and Peer Review: While the general entropy accumulation framework has been established for some time, and the initial ideas behind the MEAT were presented as conference talks in Ref. [46], publicly accessible proofs of the MEAT have only become recently available in Refs. [32, 4]. Thus, they have not yet received the adequate degree of scrutiny from the community.
  4. 4.
    Modularity: The structure of the proof is reasonably modular, with a self-contained theorem statement reducing the analysis of the overall protocol to the analysis of single rounds, in a manner that mostly does not depend on the specific details protocol.
  5. 5.
    Accessibility: We believe this proof technique to be reasonably accessible to a reader familiar with quantum information theory and numerical methods for computing QKD key rates. However, it does require the reader to familiarize themselves with a substantial amount of technical machinery.
  6. 6.
    Protocol Choices: The technique applies to a wide range of protocols. It straightforwardly accommodates noisy preprocessing, though it is not directly compatible with advantage distillation. The theoretical framework itself can handle on-the-fly announcements, but more work is needed before key rates can be computed for such scenarios.
  7. 7.
    Computability of Key Rates: The numerical techniques [50] for computing key rates under this approach appear to be somewhat harder as compared to those for the postselection technique.
8 Conclusion
This review provides a detailed and structured overview of modern proof techniques used in QKD security analysis, with a particular focus on their application to the decoy-state BB84 protocol. By identifying common gaps, misconceptions, and technical challenges in the literature, it clarifies the conceptual foundations and tools required for rigorous and practically relevant QKD security proofs.
Our aim is to serve both as a primer for newcomers entering the field and as a resource for experienced researchers seeking to critically evaluate or build upon existing proof strategies. To our knowledge, this is the most comprehensive review of QKD security proofs to date.
As demonstrated throughout this review, no single publication currently offers a complete and rigorous security proof for a fully specified, practically relevant QKD protocol.  [^63] Although, as we stated earlier, Ref. [71] is a recent promising attempt at meeting this high standard, and is currently undergoing scrutiny from the community. While significant progress has been made and most of the essential tools now exist, they have yet to be combined in a precise and correct manner to yield a certifiable proof. Achieving such a proof is critical for advancing QKD from experimental demonstrations to practical deployment. This remains an important open problem for the field.
@@ Acknowledgments
The writing of this document greatly benefited from interactions with the wider community, which includes numerous insights, discussions, and feedback, over many years. However, any errors or omissions are entirely our own responsibility. We would like to thank Takaya Matsuura for teaching the details of the phase error correction proof approach. We would like to thank Masato Koashi for further explaining the applications of the phase error correction proof approach, and the source-replacement scheme as applied to this proof technique. We thank Fred Fung for explaining the universal squasher to us. We thank Toyohiro Tsurumaru, Victor Zapatero, Margarida Pereira and Guillermo Currás-Lorenzo for helpful discussions about the EUR and phase error correction techniques, and the similarities between them. We also thank Ian George, Jie Lin and Peter Brown for helpful discussions on the EAT and its improved versions. We thank Christoph Pacher for providing detailed feedback on the first version of this work.